probability a random person will have no healur pIOD 2. A circle with a diameter equal to 18 probability of a dart hitting the square if it hits the circle. cm has a square with a side of 4 cm within it. Find the
Answer #1

Area of the circle = π (18/2)^2 = 81 π cm^2

Area of the square = 4^2 = 16 cm^2

P(Square | Circle) = P(Square and Circle)/P(Circle) = P(square)/P(Circle)

= 16/(81 π) = 56/891